OpenStudy (anonymous):

x-4/2=-5 please solve and show each step

OpenStudy (anonymous):

multiply both sides by 2 => (x-4) = -5; add four to both sides =>x = -1

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

i dont understand can you show it

OpenStudy (anonymous):

x=4/2 -5 x=(4-10)/2 x=-6/2 x=-3